OpenStudy (anonymous):

0.8t+0.2(t-0.4)=1.75

OpenStudy (anonymous):

t=1.83

OpenStudy (anonymous):

How do you find the answer?

OpenStudy (anonymous):

0.8t + 0.2t - (0.2)(0.4) = 1.75 (0.8 + 0.2)t - 0.08= 1.75 1t= 1.75 + 0.08 t= 1.83

OpenStudy (anonymous):

Thanks :-)

OpenStudy (anonymous):

use BODMAS... leave the "t" on the LHS take the constants to RHS THEN SOLVE FOR "t"